# Possible Causes of False General Relativity Violations in Gravitational Wave Observations

### v1

@@ 4.3.1 Beyond Compact Object Mergers on Bound Orbits
Parabolic or hyperbolic scattering (Junker & Schäfer, 1992) as well as head-on collision of compact objects (Davis et al., 1971; Zerilli, 1970; Simone et al., 1995) may give rise to GW signals which may resemble that of a quasi-circular CBC close to the peak of the signal. Therefore, for relatively short-duration signals, there is a risk of confusing a compact binary merger with one of the above classes of sources, leading to biases on the source parameters and thereby affecting tests of GR. In the case of GW190521, studies have discussed the degeneracy between a precessing compact binary in quasi-circular orbit with a binary that undergoes head-on collision (Calderón Bustillo et al., 2021) and a merger of two nonspinning black holes on hyperbolic orbits (Gamba et al., 2023). It is argued that the lack of premerger features in certain precessing configurations in quasi-circular CBC may mimic a head-on collision leading to underestimation of mass parameters and overestimation of luminosity distance when a quasi-circular CBC waveform is employed for parameter estimation. Obviously, such biases will directly affect most tests of GR.
@@ 4.3.2 Black Hole Mimickers
There are various exotic compact objects that are massive and compact enough that gravitational waveforms from binaries of such objects could be close to those from a binary black hole (see, e.g., Cardoso & Pani, 2019; Maggio et al., 2021). The simplest such objects can be described by GR minimally coupled to a non-Standard Model field (e.g., an ultralight scalar field describing dark matter, Ferreira, 2021). More complicated models for such objects involve nonminimally coupled fields, where it may make more sense to treat the additional scalar field as part of the gravity sector. However, even in the case where gravity is still GR, the specifics of the waveform would still differ from that of a binary black hole in GR, and one would thus obtain a false deviation from GR when applying a test of GR based on a binary black hole waveform model. The most theoretically well-modelled such objects are boson stars (see, e.g., Liebling & Palenzuela, 2023), which are formed from a massive complex scalar or vector field, that may be self-interacting, as is necessary to obtain more compact stars (that are thus more similar to black holes)—see, e.g., Sennett et al. (2017). However, there are many other models, including quite exotic objects, like gravastars (Mazur & Mottola, 2004), which have an interior made of de Sitter space. A concrete framework for these exotic objects might require GR deviations (Mottola, 2022), but they can be described also using exotic matter within GR (e.g., Uchikata et al., 2016).
@@
For all of these cases, there will be the same matter effects on the inspiral that one finds in the PN approximation for binary neutron stars (some of which are discussed in Section 3.1.2), albeit with different values. In particular, there will be effects of nonzero tidal deformabilities (see, e.g., Uchikata et al., 2016; Cardoso et al., 2017; Sennett et al., 2017; Chakraborty et al., 2023), and the excitation of resonant modes in the objects (see, e.g., Asali et al., 2020), as well as effects from multipoles that are different from those in black holes (see, e.g., Krishnendu et al., 2017; Loutrel et al., 2022) and a lack of the relatively large GW absorption (a.k.a. tidal heating) one obtains with black holes (see, e.g., Mukherjee et al., 2022). There will also be differences in the merger-ringdown part of the signal (see, e.g., Sanchis-Gual et al., 2019; Bezares et al., 2022b; Siemonsen & East, 2023, for simulations of orbiting binary boson stars). If the merger of a binary of exotic compact objects forms an ultracompact object (i.e., an object that has a light ring outside its surface), then the ringdown is nearly indistinguishable from that of a black hole and a train of modulated pulses—known as GW echoes—is emitted in the late postmerger stage (Barausse et al., 2014; Cardoso et al., 2016b). From the analysis of current GW events, no evidence for postmerger echoes has been found with unmodelled and modelled searches (Westerweck et al., 2018; Nielsen et al., 2019; Lo et al., 2019; Uchikata et al., 2019; Tsang et al., 2020; Abbott et al., 2021a, b; Miani et al., 2023), despite claims of echo detections in Abedi et al. (2017); Conklin et al. (2018); Abedi & Afshordi (2019); Abedi et al. (2023). Moreover, for perfectly reflecting objects the presence of echoes is disfavored by the current upper bounds on the stochastic background in the advanced LIGO frequency band (Barausse et al., 2018).
@@
If one has a single population of exotic stars that are formed from a single fundamental field, then the non-GR effects in the inspiral will be solely determined by the masses of the objects, and there will be a maximum mass of stable stars, just as in the neutron star case. Thus, if one can measure these effects (and the masses of the stars) accurately (using, e.g., a more refined version of the analysis given in Johnson-McDaniel et al., 2020), then one can check if the signals are indeed consistent with coming from a population of binaries of such stars. While alternate theories of gravity with an intrinsic scale will have a roughly similar behavior, where the GR deviation decreases with increasing mass of the black holes, it seems unlikely that an alternative theory of gravity would be able to mimic the situation of exotic stars to a high degree of accuracy. Moreover, if there is a population of exotic binaries as well as binary black holes, then one may observe binary black holes with very similar masses, spins, and distances as the exotic binaries, where a modified theory would predict that one would also observe deviations for the black hole binaries. Thus, while it is likely that the two situations could be confused with initial observations, it should be straightforward to distinguish them with high-accuracy observations. However, the ability of a given set of observations to distinguish specific exotic star models and specific alternative theories would need to be tested with explicit calculations.
@@
For instance, black holes can have nonzero tidal deformabilities in certain alternative theories, such as those that introduce higher-order-in-curvature corrections in the action (Cardoso et al., 2017, 2018). However, in such models the dimensionless tidal deformabilities are proportional to inverse powers of the black hole mass, $1/M^{n}$, where $n$ is a positive integer that depends on the theory ($n=4$ or $6$ in the calculations cited). This is not a good match for the mass dependence of any of the boson star models considered in Cardoso et al. (2017), and while it might be possible to find an exotic star model that gives a better match, the stars would still have a maximum mass, while the black holes in the alternative theory have nonzero tidal deformabilities for all masses. The black holes also have differences in the spin-induced multipoles (see, e.g., Cano et al., 2022) that would also have to be reproduced by the exotic stars, which is unlikely to be possible to more than moderate accuracy. For instance, for some families of boson stars, the spin-induced moments have minimum values larger than their Kerr values (similar to the minimum values of tidal deformability), and show a different spin dependence than one obtains for alternative theories (see, e.g., Vaglio et al., 2022). Additionally, there will be differences in the GW absorption comparing black holes in this theory and black hole mimickers with no horizon (which will generally have a much smaller GW absorption cross section than black holes). However, one also expects that the GW absorption in such theories will differ from that in GR due to the differences in the static tidal response, given the relation between this and GW absorption/tidal heating (see, e.g., Bhatt et al., 2023). Moreover, there are also changes to the binary’s dynamics that do not come from finite size effects in such theories (see, e.g., Sennett et al., 2020), albeit only occurring at high PN orders.
@@
In the scenario where one or both of the black holes have boson clouds around them, superradiance (see, e.g., Brito et al., 2020) will give deviations from a vacuum binary black hole signal that are similar to those that one obtains in the case of exotic compact objects. However, the same general arguments hold for distinguishing such a binary from a binary black hole in an alternative theory of gravity. Of course, in the case of boson clouds, there will not be a maximum mass of the binary’s components, and the absorption of GWs will be very similar to that of vacuum black holes. However, there will also be time dependence of the tidal deformability and non-black hole multipole moments due to perturbations or even disruption of the clouds due to the effects of the other black hole (see, e.g., Cardoso et al., 2020; Baumann et al., 2022; De Luca et al., 2023). Additionally, since the superradiant growth of the clouds is only possible for certain pairs of black hole masses and spins (see, e.g., Arvanitaki et al., 2017), this case should be easy to distinguish from the case of exotic compact objects when considering the population. Additionally, one can obtain constraints on the boson mass from the contributions from the superradiant instability to the stochastic background of GWs (Brito et al., 2017a, b).
Additionally, boson clouds are expected to emit a nearly periodic and long-duration GW signal (Brito et al., 2017a, b). No evidence of such signals is found in current GW data, which provides constraints on the ultra-light scalar boson field mass (see, e.g., Palomba et al., 2019; Sun et al., 2020; Dergachev & Papa, 2021; Abbott et al., 2022).
@@ 4.4 Statistical Assumptions of Astrophysical Population
Biases due to waveform modelling systematics can pile up when stacking multiple events in a catalog. Several studies (Gair & Moore, 2015; Moore et al., 2021; Hu & Veitch, 2023; Saini et al., 2023) show that even if systematics are under control at the level of the individual events, the accumulation of biases in a population analysis can produce false deviations from GR if the catalog is large enough. Depending on the actual population of resolved signals and on the way the events are combined, false deviations can appear with as little as $\sim 30$ events with SNR$>20$ in the most pessimistic scenarios (Moore et al., 2021). Moreover, restricting the study to golden events with high SNR is even more vulnerable to false deviations once these events become routine in XG detectors (Hu & Veitch, 2023; Saini et al., 2023), although techniques to mitigate the biases have been proposed (Gair & Moore, 2015).
@@
Furthermore, combining events requires concrete assumptions about the impact of the astrophysical population and the detectability of GW sources that violate GR. Many parameterized tests of GR infer the presence of expected correlations between individual source parameters (such as the total mass of a binary black hole system) and the deviation parameter (Psaltis et al., 2021). These correlated features within the inferred posterior distributions for individual events imply that specific choices regarding the astrophysical population distribution can skew these results to different regions of the parameter space.
@@
In a recent study, Payne et al. (2023) demonstrate that neglecting the astrophysical population leads to inferences which are $\sim 0.4\sigma$ less consistent with GR within GWTC-3 for parameterized tests of GR. However, they show that such biases can be mitigated by jointly inferring the astrophysical population properties while combining the distributions of GR violation parameters. Furthermore, Magee et al. (2024) illustrate that neglecting the loss in detectability of signals with GR violations places constraints on PN deviations that are up to $10\%$ too narrow when ignoring the selection bias in the population. These studies highlight the need to carefully consider the underlying statistical methodologies used when attempting to test GR. In the same vein, astrophysical inaccuracies or biases in the properties of a source population (e.g., imperfect mass distributions) could also lead to false GR deviations. For example, this can happen if events are detected in regions of the parameter space disfavored by astrophysical population models.
@@
Combining events to test GR also requires assumptions on the GR deviations that are being tested. If the GR modification is common among all the events (as in the case of, e.g., a nonzero graviton mass or a nonzero time variation $\dot{G}$ of Newton’s constant), one can multiply the individual, marginalized likelihoods on the deviation parameter to obtain the combined likelihood for the catalog (Zimmerman et al., 2019; Perkins et al., 2021b; Moore et al., 2021; Hu & Veitch, 2023). On the other hand, if the GR deviations are independent for each event (as may be the case if black holes have “hair”), one can multiply the individual Bayes factors in favor of GR to obtain the total evidence from the catalog (Zimmerman et al., 2019; Moore et al., 2021; Hu & Veitch, 2023). In a more general framework where the distribution of GR deviations across the catalog is a known function of the event parameters (such as masses, spins, and compactness), one would need to perform a full Bayesian hierarchical inference on the population (Zimmerman et al., 2019; Isi et al., 2019a).
@@
Studies have shown that testing GR at the population level under one of the three assumptions listed above (that all events share the same beyond-GR parameter; that modified theories introduce a new unrelated parameter for each detection; or that GR deviations across the catalog are a known function of the event parameters) can lead to the wrong conclusions if the underlying GR deviation does not satisfy the assumption (Zimmerman et al., 2019; Isi et al., 2019a). Moreover, the accumulation of biases across the catalog due to waveform systematics can change significantly depending on which method is chosen to combine multiple events (Moore et al., 2021; Hu & Veitch, 2023). Recent work by Isi et al. (2022) suggests that performing a full Bayesian analysis should be the most robust approach, but it still requires assumptions that can make the inference inherently model-dependent (Zimmerman et al., 2019).
@@
As shown by Pacilio et al. (2023), the finite size of the observed catalog will produce cosmic-variance effects that can cause to incorrectly infer deviations from GR, but a bootstrapping technique can be used to mitigate this effect.
@@ 5.1 Noise Systematics
Non-stationary and non-Gaussian noise artifacts are an ever-present analysis burden in the current generation of observatories as discussed in Sections 2.1 and 2.2. While the extent to which these artifacts will alter with upgrades to current observatories or persist in future-generation observatories remains uncertain, it is difficult to imagine that they will subside to any degree. It therefore behooves analysts to understand and mitigate these noise sources as post-processing steps before any claim of a GR violation. On the other hand, the effect of contamination from overlapping signals, whether they be super- or sub-threshold to detection, will only increase and get worse as the sensitivity of instruments gets better.
@@
For current-generation detectors, data gaps are not expected to be a problem for tests of GR because of the expected length of signals in the band and the likelihood of data gaps at precisely those times. For XG observatories, however, data gaps could become more problematic, as the signal duration increases to many hours to days, and the likelihood of gaps increases.
@@
For the current generation of observatories, uncertainties due to detector calibration do not introduce biases in parameter estimation when assuming general-relativistic waveforms, and therefore are not expected to introduce problems in tests of GR (e.g., Vitale et al., 2012, and see Section 2.5). For XG observatories, assuming an $\approx 1\%$ relative error on the amplitude, and $\approx 1^{\circ}$ error in phase, detector calibration error leads to mismatch errors of approximately $10^{-5}$, which may be problematic for tests of GR (Pürrer & Haster, 2020). Of course, this is only a dominant source of uncertainty if other sources (e.g., waveform systematics) can be mitigated below this level.

### v2

@@ 4.3.1 Beyond Compact Object Mergers on Bound Orbits
Parabolic or hyperbolic scattering [447] as well as head-on collision of compact objects [448, 449, 450] may give rise to GW signals which may resemble that of a quasi-circular CBC close to the peak of the signal. Therefore, for relatively short-duration signals, there is a risk of confusing a compact binary merger with one of the above classes of sources, leading to biases on the source parameters and thereby affecting tests of GR. In the case of GW190521, studies have discussed the degeneracy between a precessing compact binary in quasi-circular orbit with a binary that undergoes head-on collision [451] and a merger of two nonspinning black holes on hyperbolic orbits [452]. It is argued that the lack of premerger features in certain precessing configurations in quasi-circular CBC may mimic a head-on collision leading to underestimation of mass parameters and overestimation of luminosity distance when a quasi-circular CBC waveform is employed for parameter estimation. Obviously, such biases will directly affect most tests of GR.
@@ 4.3.2 Black Hole Mimickers
There are various exotic compact objects that are massive and compact enough that gravitational waveforms from binaries of such objects could be close to those from a binary black hole (see, e.g., [453, 454]). The simplest such objects can be described by GR minimally coupled to a non-Standard Model field (e.g., an ultralight scalar field describing dark matter [455]). More complicated models for such objects involve nonminimally coupled fields, where it may make more sense to treat the additional scalar field as part of the gravity sector. However, even in the case where gravity is still GR, the specifics of the waveform would still differ from that of a binary black hole in GR, and one would thus obtain a false deviation from GR when applying a test of GR based on a binary black hole waveform model. The most theoretically well-modelled such objects are boson stars (see, e.g., [456]), which are formed from a massive complex scalar or vector field, that may be self-interacting, as is necessary to obtain more compact stars (that are thus more similar to black holes)—see, e.g., [457]. However, there are many other models, including quite exotic objects, like gravastars [458], which have an interior made of de Sitter space. A concrete framework for these exotic objects might require GR deviations [459], but they can be described also using exotic matter within GR (e.g., [460]).
@@
For all of these cases, there will be the same matter effects on the inspiral that one finds in the PN approximation for binary neutron stars (some of which are discussed in Section 3.1.2), albeit with different values. In particular, there will be effects of nonzero tidal deformabilities (see, e.g., [460, 461, 457, 462]), tidal disruption (at least for sufficiently unequal-mass binaries; see, e.g., [463]), and the excitation of resonant modes in the objects (see, e.g., [464]), as well as effects from multipoles that are different from those in black holes (see, e.g., [173, 465]) and possibly a lack of the relatively large GW absorption (a.k.a. tidal heating) one obtains with black holes (see, e.g., [466]). However, since recent studies [206, 467] have shown that neutron stars can have larger GW absorption than black holes if they have a sufficiently large shear viscosity, it is possible that the same is true for some potential black hole mimickers, though this is not the case for, e.g., standard models of boson stars. There will also be differences in the merger-ringdown part of the signal (see, e.g., for simulations of orbiting binary boson stars [468, 469, 470, 471]). If the merger of a binary of exotic compact objects forms an ultracompact object (i.e., an object that has a light ring outside its surface), then the ringdown is nearly indistinguishable from that of a black hole and a train of modulated pulses—known as GW echoes—is emitted in the late postmerger stage [32, 472]. From the analysis of current GW events, no evidence for postmerger echoes has been found with unmodelled and modelled searches [473, 474, 475, 476, 477, 4, 5, 478], despite claims of echo detections in [479, 480, 481, 482]. Moreover, for perfectly reflecting objects the presence of echoes is disfavored by the current upper bounds on the stochastic background in the advanced LIGO frequency band [483].
@@
If one has a single population of exotic stars that are formed from a single fundamental field, then the non-GR effects in the inspiral will be solely determined by the masses of the objects, and there will be a maximum mass of stable stars, just as in the neutron star case. Thus, if one can measure these effects (and the masses of the stars) accurately (using, e.g., a more refined version of the analysis given in [484]), then one can check if the signals are indeed consistent with coming from a population of binaries of such stars. While alternate theories of gravity with an intrinsic scale will have a roughly similar behavior, where the GR deviation decreases with increasing mass of the black holes, it seems unlikely that an alternative theory of gravity would be able to mimic the situation of exotic stars to a high degree of accuracy. Moreover, if there is a population of exotic binaries as well as binary black holes, then one may observe binary black holes with very similar masses, spins, and distances as the exotic binaries, where a modified theory would predict that one would also observe deviations for the black hole binaries. Thus, while it is likely that the two situations could be confused with initial observations, it should be straightforward to distinguish them with high-accuracy observations. However, the ability of a given set of observations to distinguish specific exotic star models and specific alternative theories would need to be tested with explicit calculations.
@@
For instance, black holes can have nonzero tidal deformabilities in certain alternative theories, such as those that introduce higher-order-in-curvature corrections in the action [461, 485]. However, in such models the dimensionless tidal deformabilities are proportional to inverse powers of the black hole mass, $1/M^{n}$, where $n$ is a positive integer that depends on the theory ($n=4$ or $6$ in the calculations cited). This is not a good match for the mass dependence of any of the boson star models considered in [461], and while it might be possible to find an exotic star model that gives a better match, the stars would still have a maximum mass, while the black holes in the alternative theory have nonzero tidal deformabilities for all masses. The black holes also have differences in the spin-induced multipoles (see, e.g., [486]) that would also have to be reproduced by the exotic stars, which is unlikely to be possible to more than moderate accuracy. For instance, for some families of boson stars, the spin-induced moments have minimum values larger than their Kerr values (similar to the minimum values of tidal deformability), and show a different spin dependence than one obtains for alternative theories (see, e.g., [487]). Additionally, there will be differences in the GW absorption comparing black holes in this theory and black hole mimickers with no horizon (which will generally have a much smaller GW absorption cross section than black holes). However, one also expects that the GW absorption in such theories will differ from that in GR due to the differences in the static tidal response, given the relation between this and GW absorption/tidal heating (see, e.g., [488]). Moreover, there are also changes to the binary’s dynamics that do not come from finite size effects in such theories (see, e.g., [489]), albeit only occurring at high PN orders.
@@ 4.4 Statistical Assumptions of Astrophysical Population
Biases due to waveform modelling systematics can pile up when stacking multiple events in a catalog. Several studies [490, 491, 90, 136] show that even if systematics are under control at the level of the individual events, the accumulation of biases in a population analysis can produce false deviations from GR if the catalog is large enough. Depending on the actual population of resolved signals and on the way the events are combined, false deviations can appear with as little as $\sim 30$ events with SNR$>20$ in the most pessimistic scenarios [491]. Moreover, restricting the study to golden events with high SNR is even more vulnerable to false deviations once these events become routine in XG detectors [90, 136], although techniques to mitigate the biases have been proposed [490].
@@
Furthermore, combining events requires concrete assumptions about the impact of the astrophysical population and the detectability of GW sources that violate GR. Many parameterized tests of GR infer the presence of expected correlations between individual source parameters (such as the total mass of a binary black hole system) and the deviation parameter [492]. These correlated features within the inferred posterior distributions for individual events imply that specific choices regarding the astrophysical population distribution can skew these results to different regions of the parameter space.
@@
In a recent study, Payne et al. [493] demonstrate that neglecting the astrophysical population leads to inferences which are $\sim 0.4\sigma$ less consistent with GR within GWTC-3 for parameterized tests of GR. However, they show that such biases can be mitigated by jointly inferring the astrophysical population properties while combining the distributions of GR violation parameters. Furthermore, Magee et al. [494] illustrate that neglecting the loss in detectability of signals with GR violations places constraints on PN deviations that are up to $10\%$ too narrow when ignoring the selection bias in the population. These studies highlight the need to carefully consider the underlying statistical methodologies used when attempting to test GR. In the same vein, astrophysical inaccuracies or biases in the properties of a source population (e.g., imperfect mass distributions) could also lead to false GR deviations. For example, this can happen if events are detected in regions of the parameter space disfavored by astrophysical population models, such as hierarchical mergers from black holes residing in the theorized upper-mass gap [495]. Tests of GR will need to adopt the ever-growing astrophysical population knowledge to remain sufficiently unbiased.
@@
Combining events to test GR also requires assumptions on the GR deviations that are being tested. If the GR modification is common among all the events (as in the case of, e.g., a nonzero graviton mass or a nonzero time variation $\dot{G}$ of Newton’s constant), one can multiply the individual, marginalized likelihoods on the deviation parameter to obtain the combined likelihood for the catalog [496, 497, 491, 90]. On the other hand, if the GR deviations are independent for each event (as may be the case if black holes have “hair”), one can multiply the individual Bayes factors in favor of GR to obtain the total evidence from the catalog [496, 491, 90]. In a more general framework where the distribution of GR deviations across the catalog is a known function of the event parameters (such as masses, spins, and compactness), one would need to perform a full Bayesian hierarchical inference on the population [496, 498].
@@
Studies have shown that testing GR at the population level under one of the three assumptions listed above (that all events share the same beyond-GR parameter; that modified theories introduce a new unrelated parameter for each detection; or that GR deviations across the catalog are a known function of the event parameters) can lead to the wrong conclusions if the underlying GR deviation does not satisfy the assumption [496, 498]. Moreover, the accumulation of biases across the catalog due to waveform systematics can change significantly depending on which method is chosen to combine multiple events [491, 90]. Recent work by [499] suggests that performing a full Bayesian analysis should be the most robust approach, but it still requires assumptions that can make the inference inherently model-dependent [496].
@@
As shown by [500], the finite size of the observed catalog will produce cosmic-variance effects that can cause to incorrectly infer deviations from GR, but a bootstrapping technique can be used to mitigate this effect.
@@ 5.1 Noise Systematics
Non-stationary and non-Gaussian noise artifacts are an ever-present analysis burden in the current generation of observatories as discussed in Sections 2.1 and 2.2. While the extent to which these artifacts will alter with upgrades to current observatories or persist in future-generation observatories remains uncertain, it is difficult to imagine that they will subside to any degree. It therefore behooves analysts to understand and mitigate these noise sources as post-processing steps before making any claim of a GR violation. On the other hand, the effect of contamination from overlapping signals, whether they be super- or sub-threshold to detection, will only increase and get worse as the sensitivity of instruments gets better.
@@
For current-generation detectors, data gaps are not expected to pose a problem for tests of GR due to the typically short duration of signals in the band and the low likelihood of data gaps occurring precisely during those times. For XG observatories, however, data gaps could become more problematic, as the signal duration increases to many hours to days, and the likelihood of gaps increases.
@@
For the current generation of observatories, uncertainties due to detector calibration do not introduce biases in parameter estimation when assuming general-relativistic waveforms, and therefore are not expected to introduce problems in tests of GR (e.g., [102] and see Section 2.5). For XG observatories, assuming an $\approx 1\%$ relative error on the amplitude, and $\approx 1^{\circ}$ error in phase, detector calibration error leads to mismatch errors of approximately $10^{-5}$, which may be problematic for tests of GR [501]. Of course, this is only a dominant source of uncertainty if other sources (e.g., waveform systematics) can be mitigated below this level.